OpenStudy (anonymous):

Find the difference (16x^3 + 2x^3) - 4x^3 I'm having trouble figuring out whether its 16x^3 - 2x^2. Or. 12x^3 - 2x^2. Or whether I'm completely wrong.

OpenStudy (ivettef365):

all exponents are ^3, why you have x^2

OpenStudy (debbieg):

Where would a 2nd-degree term materialize from? If these are all x-cubed terms, then they are like terms so can be combined in addition & subtraction. Just combine the coefficients.

OpenStudy (anonymous):

DebbieG. Do what with the what? I don't understand.

OpenStudy (ivettef365):

you just need to add the digits the exponent stays the same so you have 16 + 2 -4 = ?

OpenStudy (anonymous):

14x^3??

OpenStudy (ivettef365):

correct

OpenStudy (debbieg):

Combine like terms. If the variables and the powers are the same, they are like terms - so you add/subtract the coefficients. Eg: \[3x^2+5x^2=(3+5)x^2=8x^2\] But you can't add/subtract non-like terms, so eg:\[3x^2+5x^3\]can't be combined, so the expression is fully simplified. Now above you have\[(16x^3+2x^3)-4x^3\]All are 3rd-degree terms (meaning the exponent is 3), so they combine as like terms. Just add/subtract the coefficients. Nothing here will give you an x-squared term though.

OpenStudy (anonymous):

Okay I got it. Thanks